# VRSD: Rethinking Similarity and Diversity for Retrieval in Large Language Models

### v1

@@ Abstract
Vector retrieval algorithms are vital for semantic queries in the evolving landscape of Large Language Models (LLMs). Retrieving vectors that simultaneously meet criteria for both similarity and diversity significantly enhances the capabilities of LLM-based agents. Despite the widespread use of the Maximal Marginal Relevance (MMR) in retrieval scenarios with relevance and diversity requirements, fluctuations caused by variations in the parameter $\lambda$ within the MMR complicate the determination of the optimization trajectory in vector spaces, thus obscuring the direction of enhancement. Moreover, there is a lack of a robust theoretical analysis for the constraints of similarity and diversity in retrieval processes. This paper introduces a novel approach to characterizing both constraints through the relationship between the sum vector and the query vector. The proximity of these vectors addresses the similarity constraint, while necessitating that individual vectors within the sum vector divergently align with the query vector to satisfy the diversity constraint. We also formulate a new combinatorial optimization challenge, taking a selection of $k$ vectors from a set of candidates such that their sum vector maximally aligns with the query vector, a problem we demonstrate to be NP-complete. This establishes the profound difficulty of pursuing similarity and diversity simultaneously in vector retrieval and lays a theoretical groundwork for further research. Additionally, we present the heuristic algorithm Vectors Retrieval with Similarity and Diversity (VRSD) which not only has a definitive optimization goal and eschews the need for preset parameters but also offers a modest reduction in time complexity compared to MMR. Empirical validation further confirm that VRSD significantly surpasses MMR across various datasets.
@@ 1 Introduction
Vector retrieval algorithms are crucial for semantic queries and have become increasingly integral to the deployment of Large Language Models (LLMs). Effective interaction with LLMs frequently necessitates the provision of relevant or similar examples to elicit enhanced responses [17]. The introduction of Retrieval Augmented Generation (RAG) has notably advanced the capabilities in knowledge-intensive tasks [16], underscoring the growing importance of retrieval methods. Empirical evidence suggests that employing the BM25 algorithm to select examples from the training set markedly improves LLMs performance over random selection [17, 19]. Moreover, leveraging existing text embedding models for example retrieval often surpasses BM25, particularly in specific contexts [26, 31]. And the advent of Dense Retrieval, which employs dense vectors for semantic matching in latent spaces [5, 15], represents a evolution over traditional sparse retrieval methods like BM25 by utilizing the robust modeling capabilities of pre-trained language models to learn relevance functions [8]. Innovations such as the applying the dual encoder framework [11] and dynamic listwise distillation [27] have further refined the effectiveness of dense retrieval techniques. Subsequent enhancements in semantic parsing and in-context learning [24], facilitated by feedback from LLMs [29], have enabled more precise example selection and improved answer accuracy. Despite ongoing advancements in retrieval methods, the broadening application scope of LLMs necessitates retrieval approaches that balance relevance with diversity—specifically, a relevance-focused diversity rather than an unrestricted diversity. Additionally, the RAG framework’s ability to augment the LLMs’ external data access also underscores the need for simple yet efficient algorithms that can streamline the retrieval process.
@@
Considering the balance between similarity and diversity, the Maximal Marginal Relevance (MMR) [4] is an effective heuristic algorithm and has been widely applied in vector retrieval practices. Aiming to achieve an optimal balance, MMR incorporates a parameter, $\lambda$, which adjusts the weight of relevance and diversity by varying its value. Nevertheless, this method is not always effective; in different scenarios, $\lambda$ needs to take different values, which cannot be known in advance. Recent research [29, 32] has also explored using LLMs to enhance retrieval results, while also suggests considering the selection of a set of examples from a combinatorial optimization perspective, rather than selecting examples one by one, as the in-context examples can influence each other. In light of this, we propose using the sum vector to characterize both similarity and diversity in vector retrieval. Simply put, this involves maximizing the similarity between the sum vector of the selected vectors and the query vector, and maximizing the similarity of the sum vector to the query vector imposes a similarity constraint. At the same time, from a geometric perspective, the requirement for the sum vector to be similar to the query vector means that the selected vectors approach the query vector from different directions, thus imposing a diversity constraint. Additionally, the idea of considering the similarity between the sum vector and the query vector is analogous to the famous finding in word2vec (king - man + woman = queen) [22], as both involve obtaining complex semantic similarities through simple vector arithmetic. Therefore, using the sum vector to characterize similarity and diversity constraints not only considers similarity while reducing redundancy but also enhances the complementarity among retrieval results.
@@
  • We propose using the similarity between the sum vector and the query vector to characterize similarity and diversity constraints in vector retrieval. We formulate a novel optimization problem where we seek to select several vectors from a set of candidates such that the similarity between the sum vector of the selected vectors and the query vector is maximized.
@@
The remainder of this paper is organized as follows: Section 2 provides a review of related work in the field. Section 3 defines the problem under investigation, examines their computational complexities, and proposes a heuristic algorithm for addressing them. Section 4 presents an experimental evaluation of our heuristic algorithm. Finally, Section 5 concludes the paper and discusses potential directions for future research.
2 Related Work
2.1 Retrieval and Large Language Model
Retrieval methods in Large Language Models (LLMs) have gained traction, particularly due to their pivotal role in open-domain question answering, evidenced by seminal contributions in the field [5, 9, 14, 25, 10]. The introduction of Retrieval Augmented Generation (RAG) further underscored the significance of these methods across knowledge-intensive tasks [16], notably enhancing generation capabilities for open-domain queries [20]. This spurred the adoption of techniques such as K-Nearest-Neighbor (KNN) in diverse applications, ranging from the customization of multilingual models in machine translation [12] to improving the prediction of rare patterns in LLMs [13, 1]. Continued advancements in retrieval techniques have focused on identifying highly informative examples to augment in-context learning, thereby enabling LLM-based systems to achieve significant performance improvements with minimal examples [3]. Early on, traditional sparse retrieval methods like BM25 [28]—an extension of TF-IDF—were utilized to refine in-context learning [17]. Subsequently, the integration of LLMs’ intrinsic capabilities [30] and Sentence-BERT (SBERT) [26] facilitated the retrieval of highly pertinent examples for prompt integration. The advent of dense retrievers signified a methodological enhancement in retrieval from a machine learning perspective, and the incorporation of feedback signals with contrastive learning has yielded more effective retrieval systems [29, 32]. Recent innovations like UPRISE [6] and PRAC [23] have further optimized the performance of in-context learning by retrieving demonstrations directly from training data. Despite these advances, most retrieval methods still treat each candidate independently, which can lead to suboptimal outcomes due to the interaction effects among in-context examples, resulting in a lack of diversity. Given the expanding applications of LLMs, diversity becomes increasingly crucial, as complex inference tasks often require undefined example types. Incorporating a diverse range of examples enriches LLMs’ learning processes, facilitating more innovative and robust responses, especially for complex open-ended questions. Furthermore, a relevance-focused diversity retrieval method could help mitigate the impact of malicious information as LLMs continue to scale and assimilate more societal data. Even if some malicious information is inadvertently retrieved, the inclusion of diverse data within the same batch can enable LLMs to fully comprehend the query’s intent and provide accurate responses. Additionally, as LLMs are tasked with processing ever-greater volumes of information, the necessity for a straightforward and efficient retrieval method intensifies, minimizing the dependency on resource-intensive preparatory tasks.
2.2 Maximal Marginal Relevance
To enhance retrieval processes by accounting for both relevance and diversity, the Maximal Marginal Relevance (MMR) algorithm was introduced [4]. MMR addresses the balance between relevance and diversity in traditional retrieval and summarization methods by employing "marginal relevance" as an evaluation metric. This metric is defined as a linear combination of independently measured relevance and novelty, formulated as Eq.1:
@@ 3.1 Problem definition and complexity analysis
To address the problem of selecting a subset of vectors from a set of candidate vectors that satisfy both similarity and diversity requirements, we refer to the MMR algorithm and several LLM-based algorithms, typically considering the following premises: Firstly, the candidate vectors are identified from the entire set of vectors ($size=N$) using similarity metrics, resulting in a subset of vectors ($size=n$). Consequently, this set of candidate vectors inherently exhibits a relative high degree of similarity to the query vector. Secondly, within these $n$ candidate vectors, the vector most similar to the query vector is typically selected first, as is the case with the MMR algorithm and others. This approach is favored because, in applications such as in-context learning with LLMs, examples with the highest similarity to the query are generally the most helpful.
@@
Next, we will prove this decision problem is NP-complete. For the sake of concise proof, we further restrict the components of vectors to integers. The proof strategy is to reduce the subset sum problem [7] to this decision problem.
@@
$[t^{\prime}_{1},1]\ldots[t^{\prime}_{k},1]$ are the selected $k$ vectors. It implies that $\alpha=1$ and $t^{\prime}_{1}+\ldots+t^{\prime}_{k}=t$. Thus, this provides an affirmative answer to the $k$-subset sum problem instance $(T,t,k)$. Conversely, if vectors retrieval provides a negative answer, then a negative answer to the $k$-subset sum problem can also be obtained. The above reduction process can be clearly completed in polynomial time. Therefore, the decision problem of vectors retrieval is NP-complete. ∎
@@ 3.2 Heuristic algorithm for vectors retrieval
Since the vectors retrieval problem $(R,q,k)$ is a NP-complete problem, necessitating the use of heuristic methods to derive feasible solutions. Specifically, given a set of candidate vectors with high similarity, the objective is to select $k$ vectors that maximize the cosine similarity between the sum vector of the $k$ selected vectors and the query vector. We propose a new algorithm denoted as Vectors Retrieval with Similarity and Diversity (VRSD). VRSD initially selects the vector most similar to the query vector and then iteratively selects additional vectors from the remaining candidates. In each iteration, it chooses the vector that maximizes the cosine similarity between the cumulative sum of all previously selected vectors and the query vector, continuing this process until $k$ vectors are chosen. Further details about the VRSD algorithm can be found in Algorithm.1 and Fig.2.
Figure 2: An illustration of how the VRSD works for $(R,q,3)$. (1) The vector $d_{0}$ is first selected as it has the maximum cosine similarity (MCS) with the query vector q and. And $d_{0}$ will be removed from R. (2) The vector $d_{y}$ is second selected because, when added to $d_{0}$, their sum vector has the maximum cosine similarity with q compared to the other unselected vectors, and is also subsequently removed from R. (3) The vector $d_{z}$ is third selected as it has the maximum cosine similarity with q after being added to the sum vector.
@@
3.3 Time complexity analysis of VRSD algorithm
@@ 4.1 Experiments detail
  • ARC-DA [2]: A dataset of direct-answer science questions derived from the ARC multiple-choice question. Each example contains a question and multiple answers.
@@
  • OpenBookQA [21]: A dataset of multiple-choice science questions, which probe the understanding of science facts and the application of these facts to novel situations. Each example contains a question, multiple choices, and an answer.
@@
  • Puzzle [18]: A question answering dataset. These questions belong to lateral thinking puzzle. Each example contains a question and an answer.
@@
For each item in each datasets, we concatenate the question part with its corresponding answer, subsequently selecting 20% of these concatenated items to form the test set, wherein the question parts are isolated. Items designated for the test set are excluded from the original dataset for subsequent experiments, where four examples are retrieved for each test question. Our evaluation focuses on two primary aspects: retrieval quality and answer quality. Retrieval quality is assessed by aggregating four vectors retrieved using either VRSD or MMR into a sum vector—denoted as $d_{\text{VRSD}}$ and $d_{\text{MMR}}$—which reflects the vectorial direction from which the examples approach the query vector $q$. We compute the cosine similarity between the sum vectors and the query vector as $\cos(d_{\text{VRSD}},q)$ and $\cos(d_{\text{MMR}},q)$. The comparison includes counting instances where $\cos(d_{\text{VRSD}},q)$ exceeds $\cos(d_{\text{MMR}},q)$, termed as the VRSD win. rate, and calculating the maximum difference between these cosine similarities for all queries in each test set. Additionally, we compute the mean cosine similarity values for these vectors. Such metrics are instrumental in elucidating the algorithms’ capacity to balance similarity and diversity. For answer quality assessment, we reconstruct prompts by concatenating the original sentences corresponding to the four retrieved vectors with the initial question, and input these into large LLMs. The responses generated by open-source LLM open-mistral-7b and closed-source LLM gpt-3.5-turbo are then compared with standard answers using the Rogue-L metric for ARC-DA and Puzzle datasets, and Accuracy for OpenBookQA. This evaluation helps us ascertain the efficacy of the retrieved examples in providing accurate answers
4.2 Experiments Results
Table 1: Algorithm’s performance in different dataset. Max-diff displays the maximum difference between $\cos(d_{\text{VRSD}},q)$ and $\cos(d_{\text{MMR}},q)$ across all queries in the test set. Mean displays the mean value of $\cos(d_{\text{VRSD}},q)$ and $\cos(d_{\text{MMR}},q)$ in the test set. Boldface indicates the algorithm with the highest mean value under different dataset.
        Algorithm         VRSD win.rate         Max-diff         Mean
        ARC-DA         VRSD         -         -         0.740
        MMR($\lambda$=0)         97.7%         0.160         0.696
        MMR($\lambda$=0.5)         92.5%         0.108         0.720
        MMR($\lambda$=1)         95.3%         0.158         0.710
        OpenBookQA         VRSD         -         -         0.833
        MMR($\lambda$=0)         97.3%         0.135         0.809
        MMR($\lambda$=0.5)         92.6%         0.101         0.822
        MMR($\lambda$=1)         96.8%         0.102         0.812
        Puzzle         VRSD         -         -         0.592
        MMR($\lambda$=0)         100%         0.161         0.537
        MMR($\lambda$=0.5)         90%         0.052         0.576
        MMR($\lambda$=1)         100%         0.132         0.577
Table 2: Algorithm’s performance under different LLMs. Boldface indicates the algorithm with the highest score under different models.

### v2

@@ Abstract
Vector retrieval algorithms are essential for semantic queries within the rapidly evolving landscape of Large Language Models (LLMs). The ability to retrieve vectors that satisfy both similarity and diversity criteria substantially enhances the performance of LLMs. Although Maximal Marginal Relevance (MMR) is widely employed in retrieval scenarios requiring relevance and diversity, variations in the parameter $\lambda$ lead to fluctuations that complicate the optimization trajectory in vector spaces. This obscures the direction of improvement and highlights the lack of a robust theoretical analysis regarding similarity and diversity constraints in retrieval processes. To address these challenges, this paper introduces a novel approach that characterizes both constraints through the relationship between the sum vector and the query vector. The proximity of these vectors ensures the similarity constraint, while requiring individual vectors within the sum vector to diverge in their alignment with the query vector satisfies the diversity constraint. We first formulate a new combinatorial optimization problem, selecting $k$ vectors from a candidate set such that their sum vector maximally aligns with the query vector, and demonstrate that this problem is NP-complete. This result underscores the inherent difficulty of simultaneously achieving similarity and diversity in vector retrieval, thereby providing a theoretical foundation for future research. Subsequently, we present the heuristic algorithm Vectors Retrieval with Similarity and Diversity, VRSD, which features a clear optimization objective and eliminates the need for preset parameters. VRSD also achieves a modest reduction in time complexity compared to MMR. Empirical validation confirms that VRSD significantly outperforms MMR across various datasets, while also demonstrating that the sum vector effectively captures both diversity and similarity simultaneously.
@@ 1 Introduction
Vector retrieval algorithms are crucial for semantic queries and have become increasingly integral to the deployment of Large Language Models (LLMs). Effective interaction with LLMs frequently necessitates the provision of relevant or similar examples to elicit enhanced responses (Liu et al., 2022). The introduction of Retrieval Augmented Generation (RAG) has notably advanced the capabilities in knowledge-intensive tasks (Lewis et al., 2020), underscoring the growing importance of retrieval methods. Empirical evidence suggests that employing the BM25 algorithm to select examples from the training set markedly improves LLMs performance over random selection (Liu et al., 2022; Luo et al., 2023). Moreover, leveraging existing text embedding models for example retrieval often surpasses BM25, particularly in specific contexts (Reimers & Gurevych, 2019; Wang et al., 2022). And the advent of Dense Retrieval, which employs dense vectors for semantic matching in latent spaces (Chen et al., 2017; Lee et al., 2019), represents a evolution over traditional sparse retrieval methods like BM25 by utilizing the robust modeling capabilities of pre-trained language models to learn relevance functions (Devlin et al., 2019). Innovations such as the applying the dual encoder framework (Karpukhin et al., 2020) and dynamic listwise distillation (Ren et al., 2021) have further refined the effectiveness of dense retrieval techniques. Subsequent enhancements in semantic parsing and in-context learning (Pasupat et al., 2021), facilitated by feedback from LLMs (Rubin et al., 2022), have enabled more precise example selection and improved answer accuracy. Despite ongoing advancements in retrieval methods, the broadening application scope of LLMs necessitates retrieval approaches that balance relevance with diversity—specifically, a relevance-focused diversity rather than an unrestricted diversity. Additionally, the RAG framework’s ability to augment the LLMs’ external data access also underscores the need for simple yet efficient algorithms that can streamline the retrieval process.
@@
Considering the balance between similarity and diversity, the Maximal Marginal Relevance (MMR) (Carbonell & Goldstein, 1998) is an effective algorithm and has been widely applied in vector retrieval practices. Aiming to achieve an optimal balance, MMR incorporates a parameter, $\lambda$, which adjusts the weight of relevance and diversity by varying its value. Nevertheless, this method is not always effective; in different scenarios, $\lambda$ needs to take different values, which cannot be known in advance. Recent research (Rubin et al., 2022; Wang et al., 2023) has also explored using LLMs to enhance retrieval results, while also suggests considering the selection of a set of examples from a combinatorial optimization perspective, rather than selecting examples one by one, as the in-context examples can influence each other. In light of this, we propose using the sum vector to characterize both similarity and diversity in vector retrieval. Simply put, this involves maximizing the similarity between the sum vector of the selected vectors and the query vector, and maximizing the similarity of the sum vector to the query vector imposes a similarity constraint. At the same time, from a geometric perspective, the requirement for the sum vector to be similar to the query vector means that the selected vectors approach the query vector from different directions, thus imposing a diversity constraint. Additionally, the idea of considering the similarity between the sum vector and the query vector is analogous to the famous finding in word2vec (king - man + woman = queen) (Mikolov et al., 2013), as both involve obtaining complex semantic similarities through simple vector arithmetic. Therefore, using the sum vector to characterize similarity and diversity constraints not only considers similarity while reducing redundancy but also enhances the complementarity among retrieval results.
@@
  • We propose using the sum vector to characterize similarity and diversity constraints in vector retrieval. We formulate a novel optimization problem where we seek to select several vectors from a set of candidates such that the similarity between the sum vector of the selected vectors and the query vector is maximized.
@@
2 Theoretical analysis of MMR
2.1 Limitations of MMR
To enhance retrieval processes by accounting for both relevance and diversity, the Maximal Marginal Relevance (MMR) algorithm was introduced (Carbonell & Goldstein, 1998). MMR addresses the balance between relevance and diversity in traditional retrieval and summarization methods by employing "marginal relevance" as an evaluation metric. This metric is defined as a linear combination of independently measured relevance and novelty, formulated as Eq.1:
@@ 2.1 Limitations of MMR
2.2 Sum Vector for Retrieval
To address the challenges associated with MMR, we propose utilizing the sum vector of selected vectors to simultaneously capture both similarity and diversity. Figure.1 provides an intuitive geometric illustration of vector diversity and similarity. In practical applications, similarity does not necessarily imply strict alignment between vectors; rather, a small angle between vectors suffices to indicate similarity. A smaller angle between two vectors denotes a higher degree of similarity, while a larger angle signifies lower similarity. In the context of an embedded query, retrieval typically involves ranking candidate vectors based on the angles they form with the query vector, from smallest to largest, and selecting them according to the required quantity, for instance, by directly applying cosine similarity. However, if the selected batch of vectors for a query includes vectors with both small and large angles relative to the query vector, the batch can be considered to satisfy the diversity requirement, as it incorporates vectors that are less directly relevant. Consequently, the selected vectors do not merely satisfy the criterion of having the smallest possible angles with the query vector but instead exhibit a range of angles. Nevertheless, diversity in vector retrieval should be grounded in similarity and should not entirely deviate from it. In other words, diversity should not be achieved by deliberately selecting vectors that are entirely different in angle from the query vector. The selected examples should exhibit substantial differences while satisfying the similarity constraint, thereby complementing each other and fulfilling both similarity and diversity requirements. And Sum vector can achieve this goal well. The sum vector vector of the selected vectors is required to closely approximate the query vector, thereby effectively satisfying the similarity constraint. While from a geometric perspective, the sum vector’s similarity to the query vector implies that the selected vectors approach the query vector from different directions, indicating significant differences among the selected vectors while maintaining complementarity, which reflects diversity. Therefore, the sum vector effectively captures both diversity and similarity.
@@ 3.1 Problem definition and complexity analysis
To address the problem of selecting a subset of vectors from a set of candidate vectors that satisfy both similarity and diversity requirements, we refer to the MMR algorithm and several LLM-based algorithms, typically considering the following premises: Firstly, the candidate vectors are identified from the entire set of vectors ($size=N$) using cosine similarity metrics, resulting in a subset of vectors ($size=n$). Consequently, this set of candidate vectors inherently exhibits a relative high degree of similarity to the query vector. Secondly, within these $n$ candidate vectors, the vector most similar to the query vector is typically selected first, as is the case with the MMR algorithm and others. This approach is favored because, in applications such as in-context learning with LLMs, examples with the highest similarity to the query are generally the most helpful.
@@
Next, we will prove this decision problem is NP-complete. For the sake of concise proof, we further restrict the components of vectors to integers. The proof strategy is to reduce the subset sum problem Cormen et al. (2009) to this decision problem.
@@
$[t^{\prime}_{1},1]\ldots[t^{\prime}_{k},1]$ are the selected $k$ vectors. It implies that $\alpha=1$ and $t^{\prime}_{1}+\ldots+t^{\prime}_{k}=t$. Thus, this provides an affirmative answer to the $k$-subset sum problem instance $(T,t,k)$. Conversely, if vectors retrieval provides a negative answer, then a negative answer to the $k$-subset sum problem can also be obtained. The above reduction process can be clearly completed in polynomial time. Therefore, the decision problem of vectors retrieval is NP-complete.
@@ 3.2 Heuristic algorithm for vectors retrieval
Since the vectors retrieval problem $(R,q,k)$ is a NP-complete problem, necessitating the use of heuristic methods to derive feasible solutions. Specifically, given a set of candidate vectors with high similarity, the objective is to select $k$ vectors that maximize the cosine similarity between the sum vector of the $k$ selected vectors and the query vector. We propose a new algorithm denoted as Vectors Retrieval with Similarity and Diversity (VRSD). VRSD initially selects the vector most similar to the query vector and then iteratively selects additional vectors from the remaining candidates. In each iteration, it chooses the vector that maximizes the cosine similarity between the cumulative sum of all previously selected vectors and the query vector, continuing this process until $k$ vectors are chosen. Further details about the VRSD algorithm can be found in Algorithm.1.
@@
3.3 Time complexity analysis of VRSD
@@ 4.1 Implementation Details
  • ARC-DA (Bhakthavatsalam et al., 2021): A dataset of direct-answer science questions derived from the ARC multiple-choice question. Each example contains a question and multiple answers.
@@
  • OpenBookQA (Mihaylov et al., 2018): A dataset of multiple-choice science questions, which probe the understanding of science facts and the application of these facts to novel situations. Each example contains a question, multiple choices, and an answer.
@@
  • Puzzle (Liu et al., 2023): A question answering dataset. These questions belong to lateral thinking puzzle. Each example contains a question and an answer.
For each item in each datasets, we concatenate the question part with its corresponding answer, subsequently selecting 20% of these concatenated items to form the test set, wherein the question parts are isolated. Items designated for the test set are excluded from the original dataset for subsequent experiments, where same amount of examples are retrieved for each test question.
@@
As extensively discussed in Section 2, sum vector can well capture the similarity and diversity simultaneously. Consequently, retrieval quality is assessed by aggregating all vectors retrieved using either VRSD or MMR into a sum vector—denoted as $d_{\text{VRSD}}$ and $d_{\text{MMR}}$—which reflects the vectorial direction from which the examples approach the query vector $q$. We compute the cosine similarity between the sum vectors and the query vector as $\cos(d_{\text{VRSD}},q)$ and $\cos(d_{\text{MMR}},q)$. The comparison includes counting percentage where $\cos(d_{\text{VRSD}},q)$ exceeds $\cos(d_{\text{MMR}},q)$, termed as the VRSD win. rate, and calculating the maximum difference (Max-diff) between these cosine similarities for all queries in each test set. Additionally, we compute the mean cosine similarity values (Mean) for these vectors under each methods, respectively. Such metrics are instrumental in elucidating the algorithms’ capacity to balance similarity and diversity. The comparison was conducted on one open source model all-mpnet-base-v2 and two close source models text-embedding-3-small&text-embedding-ada-002.
The author of MMR conducted a manual evaluation of the retrieved examples, noting that ’users were asked to extract information from documents without being informed about the order in which the documents were presented—only that either "method R" or "method S" was applied.’ (Carbonell & Goldstein, 1998). Manual evaluation, however, is time-consuming and labor-intensive. We propose utilizing the sum vector for evaluation, which offers a more direct and efficient approach. As discussed earlier, the sum vector can simultaneously account for both diversity and similarity.
4.2 Experimental Results
4.2.1 Retrieval performance
Table 1: Results of MMR and VRSD on three datasets. The best results are bolded.
Model Algorithm ARC-DA OpenBookQA Puzzle
VRSD win.rate Max-diff Mean VRSD win.rate Max-diff Mean VRSD win.rate Max-diff Mean
all-mpnet-base-v2 VRSD - - 0.740 - - 0.833 - - 0.592
MMR($\lambda$=0) 97.7% 0.160 0.696 97.3% 0.135 0.809 100% 0.161 0.537
MMR($\lambda$=0.5) 92.5% 0.108 0.720 92.6% 0.101 0.822 90% 0.052 0.576
MMR($\lambda$=1) 95.3% 0.158 0.710 96.8% 0.102 0.812 100% 0.132 0.577
text-embedding-3-small VRSD - - 0.695 - - 0.751 - - 0.572
MMR($\lambda$=0) 98.8% 0.145 0.652 98.1% 0.137 0.724 100% 0.184 0.555
MMR($\lambda$=0.5) 92.3% 0.113 0.676 93.3% 0.130 0.738 90% 0.052 0.527
MMR($\lambda$=1) 96.5% 0.129 0.669 97.3% 0.085 0.729 100% 0.132 0.518
text-embedding-ada-002 VRSD - - 0.905 - - 0.922 - - 0.893
MMR($\lambda$=0) 98.5% 0.040 0.895 97.2% 0.043 0.915 100% 0.025 0.881
MMR($\lambda$=0.5) 90.8% 0.027 0.901 88.1% 0.032 0.918 95% 0.020 0.888
MMR($\lambda$=1) 95.0% 0.041 0.897 95.0% 0.028 0.915 95% 0.018 0.887
Table 2: The performance of retrieved examples under each method with different LLMs. We calculate the standard error of the mean of each value. The best results are bolded.